Use a table of values and 3. Graph the function f(x) = 2 - 2 plot at least 3 points.
Answer #1

fLa a a 90 이 Q_ 2. 2 \_ 3-1 그O a 일e 22 = 4-2= 2 늘 9 - 일 니 쟁 22) Lliol ㅋa LO) -1 ) 네를